​Discount-Mart, a major East Coast​ retailer, wants to determine the economic order quantity for its halogen lamps. It currently buys all halogen lamps from Specialty Lighting​ Manufacturers, in Atlanta. Annual demand is 2 comma 000 ​lamps, ordering cost per order is ​$0.50​, carrying cost per lamp is ​$20. As part of its new JIT​ program, Discount-Mart has signed a​ long-term contract with Specialty Lighting and will place orders electronically for its halogen lamps. ​a) What is the economic order​ quantity? nothing units per order ​(round your response to the nearest whole​ number). ​b) How many orders will be​ placed? nothing per year ​(round your response to the nearest whole​ number). ​c) What is the total annual cost of managing the inventory with this​ policy? ​$ nothing ​(round your response to the nearest whole​ number).
Answer #1

1.

EOQ = SQRT(2 * D * S / H)
EOQ = SQRT(2 * 2000 * 0.5 / 20)
EOQ = 10


2.

ORDER PER YEAR = DEMAND / EOQ = 2000 / 10 = 200

3.

TOTAL COST OF MANAGING THE INVENTORY = ANNUAL HOLDING COST + ANNUAL ORDERING COST

ANNUAL HOLDING COST = (EOQ / 2) * HOLDING COST = (10 / 2) * 20 = 100
ANNUAL ORDERING COST = ORDER PER YEAR * ORDERING COST = 200 * 0.5 = 100

TOTAL COST OF MANAGING = 100 + 100 = 200

~~~@#PLEASE LEAVE A THUMBS UP FOR THE ANSWER@~~~